# GraLoRA: Granular Low-Rank Adaptation for Parameter-Efficient Fine-Tuning

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) is a popular method for parameter-efficient fine-tuning (PEFT) of generative models, valued for its simplicity and effectiveness. Despite recent enhancements, LoRA still suffers from a fundamental limitation: overfitting when the bottleneck is widened. It performs best at ranks 32–64, yet its accuracy stagnates or declines at higher ranks, still falling short of full fine-tuning (FFT) performance. We identify the root cause as LoRA’s structural bottleneck, which introduces gradient entanglement to the unrelated input channels and distorts gradient propagation. To address this, we introduce a novel structure, Granular Low-Rank Adaptation (GraLoRA) that partitions weight matrices into sub-blocks, each with its own low-rank adapter. With negligible computational or storage cost, GraLoRA overcomes LoRA’s limitations, effectively increases the representational capacity, and more closely approximates FFT behavior. Experiments on code generation and commonsense reasoning benchmarks show that GraLoRA consistently outperforms LoRA and other baselines, achieving up to +8.5% absolute gain in Pass@1 on HumanEval+. These improvements hold across model sizes and rank settings, making GraLoRA a scalable and robust solution for PEFT. Code, data, and scripts are available at https://github.com/SqueezeBits/GraLoRA.git
@@ 1 Introduction
Task-specific fine-tuning enables a wide range of applications and significantly improves the quality and effectiveness of generative models. However, the massive scale of these models poses substantial challenges for practical deployment. To address these limitations, Parameter-Efficient Fine-Tuning (PEFT) methods have emerged as a cost-effective alternative [10, 26]. Among them, Low-Rank Adaptation (LoRA) [11] has gained particular attention for its simplicity and effectiveness, introducing trainable low-rank matrices while keeping the pre-trained model weights frozen. Although the imposed rank-$r$ bottleneck may lead to slight performance degradation compared to full fine-tuning (FFT), its efficiency has led to widespread adoption in practice.
To maximize the benefits of LoRA, various studies have proposed techniques such as improved initialization [5, 18, 20, 23] and structural refinements [9, 13, 14, 15] to enhance fine-tuning quality. While these efforts have advanced performance, a substantial quality gap remains compared to FFT, largely due to the inherent upper bound on the rank. Although using a higher rank, within hardware limits, appears to be a natural solution, unfortunately, current implementations of LoRA and its variants do not support such flexibility. Simply increasing the rank often leads to degraded accuracy in many scenarios.
@@ 2.1 Introduction to LoRA
$\frac{\partial L}{\partial R}=\frac{\partial L}{\partial Y}X^{\top}\in\mathbb{R}^{M\times N}$ (2)
@@
These result in the following update in the fused weight space:
$\frac{\partial L}{\partial R}=\frac{\partial L}{\partial B}A^{\top}+B\frac{\partial L}{\partial A^{\top}}=\frac{\partial L}{\partial Y}X^{\top}AA^{\top}+BB^{\top}\frac{\partial L}{\partial Y}X^{\top}.$ (4)
@@ 2.2 Why Does LoRA Suffer from a Larger Rank?
These findings reveal a fundamental misalignment between LoRA updates and the gradient landscape shaped by FFT. The entangled influence of input channels caused by the low-rank projection limits LoRA’s ability to selectively learn from salient features, particularly under skewed input statistics. While the negative impact of outliers has been well recognized in the context of quantization-aware training [25] [16], their influence on LoRA’s behavior has not been systematically studied until now.
@@ 3.4 Tradeoff Analysis
Selection of k While GraLoRA increases the total rank from $r$ to $kr$, each individual block, represented as $B_{i,j}A_{i,j}^{\top}\in\mathbb{R}^{\frac{M}{k}\times\frac{N}{k}}$, is constrained to a reduced rank of $\frac{r}{k}$. As a result, increasing $k$ beyond a certain threshold can degrade performance due to limited expressiveness within each block. This effect is especially pronounced when the overall rank $r$ is small. Empirically, we observed that maintaining a minimum block expressiveness of approximately $r/k^{2}\approx 8$ yields stable performance across various configurations. Based on this observation, we adopted $k=2$ for ranks 16 and 32, and $k=4$ for ranks 64 and 128 in our experiments. Detailed results from the $k$-sweep can be found in Section 4.4.
@@ 4 Experiments
In order to validate the superiority of the proposed idea, we conduct an extensive analysis on large-scale dataset with the state-of-the art LLMs. We evaluate GraLoRA across two challenging domains: code generation and commonsense reasoning. Our experiments are designed to assess whether the proposed granular adaptation mechanism improves performance across varying model sizes, LoRA ranks, and tasks that require nuanced reasoning and high representational fidelity.
@@ 4.1 Experimental Setup
Code Generation. We fine-tuned each model on the Magicoder-Evol-Instruct-110k [24] train dataset, a curated and decontaminated subset of WizardCoder [17], comprising high-quality instruction–response pairs for programming tasks. Evaluation was conducted on the Humaneval+ test dataset following He et al. [9], which samples 50 completions per problem using a temperature of 0.2. We report Pass@1, Pass@5, and Pass@10 accuracy following standard protocol via BigCode Evaluation Harness [1]. We have evaluated the last epoch weight across all methods.
Commonsense Reasoning We fine-tune each model across 8 commonsense tasks: BoolQ [6], PIQA [4], SIQA [22], HellaSwag [28], WinoGrande [21], ARC-Challenge, ARC-Easy [7], and OpenBookQA [19]. We followed the training pipeline proposed by LLM-Adapters [12], using a merged dataset composed of the training sets from all tasks. Evaluation was conducted on individual testing dataset for each task on the last epoch weight. Detailed training parameters can be found in Appendix D.
Training Details We conducted experiments on four open-sourced LLMs with different architecture and size: LLaMA3.1–8B, LLaMA3.1–70B ( [8]), Qwen-2.5-1.5B, and Qwen-2.5-7B ( [27]). We used pre-trained models rather than instruction-tuned models following the common practice ( [14, 15]). We applied PEFT methods on all linear modules from attention ( $W_{q},W_{k},W_{v},W_{o}$ )and feed-forward networks ( $W_{up},W_{down},W_{gate}$). We set the hyper-parameters based on the optimal configurations from Biderman et al.  [3] and He et al. [9], employing decoupled LionW optimizer with batch size of 192 and setting LoRA $\alpha=2r$. We applied alpaca-chat template for both tasks. Code generation was conducted on 4 A100 80G GPUs and commonsense reasoning task was conducted on 2 H100 80G GPUs for 1.5 8B models. 70B model was conducted on 8 A100 80G GPUs. We compared GraLoRA to three representative PEFT methods: LoRA, MoRA  [14] and RaSA  [9].
@@ 4.2 Results on Code Generation
16 LoRA 6.2h 1.00$\times$ 56.1% 65.3% 68.1%
MoRA 8.8h 1.42$\times$ 53.6% 62.2% 64.5%
RaSA 6.7h 1.08$\times$ 53.7% 64.4% 66.7%
GraLoRA* 6.7h 1.08$\times$ 58.0% 67.1% 70.1%
32 LoRA 6.5h 1.00$\times$ 58.4% 68.0% 69.9%
MoRA 9.1h 1.40$\times$ 58.3% 66.7% 69.0%
RaSA 6.8h 1.05$\times$ 57.2% 67.9% 70.5%
GraLoRA 6.9h 1.06$\times$ 58.9% 67.0% 69.0%
64 LoRA 6.7h 1.00$\times$ 58.1% 66.4% 68.5%
MoRA 9.7h 1.45$\times$ 57.2% 66.4% 69.2%
RaSA 6.9h 1.03$\times$ 56.6% 65.4% 67.9%
GraLoRA 7.2h 1.07$\times$ 60.5% 71.2% 72.6%
128 LoRA 7.0h 1.00$\times$ 55.8% 64.8% 68.6%
MoRA 9.9h 1.41$\times$ 52.8% 62.3% 65.3%
RaSA 7.6h 1.09$\times$ 57.5% 65.5% 67.5%
GraLoRA 7.7h 1.10$\times$ 64.3% 71.7% 73.7%
@@ 4.3 Results on Commonsense Reasoning
Table 3: Commonsense reasoning accuracy across models and tasks. All values are percentages; bold indicates the best performance per row. HS means HellaSwag, and WG WinoGrande.
@@
As shown in Table 3, GraLoRA outperformed other methods across a wide range of models and tasks. Notably, GraLoRA demonstrated superior performance across models of varying scales, achieving a 1.1% improvement in average accuracy on both Qwen2.5-1.5B and LLaMA3.1-70B. It also delivered a 0.9% gain on the widely used mid-sized model, Qwen2.5-7B.
Furthermore, GraLoRA achieved the best results on 20 out of 24 tasks, consistently outperforming alternatives across benchmarks. These results support our analysis in Section 3.3, showing that GraLoRA’s localized updates enhance alignment with FFT and promote robust generalization in multi-aspect reasoning tasks.
@@ 4.4 Ablation Study
Figure 8: (a) GraLoRA $k$ sweep results and (b) Hybrid GraLoRA Ratio sweep results for LLaMA3.1–8B on code generation task. Ratio 0 implies default GraLoRA and ratio 1 implies vanilla LoRA in (b).
@@ 5 Conclusion
GraLoRA addresses this limitation by dividing the adaptation space into $k^{2}$ independently trained low-rank adapters, enabling spatially localized and context-aware updates. Our theoretical analysis shows that this design increases expressivity by a factor of $k$, without additional parameters or computational cost. Moreover, under outlier activations, GraLoRA effectively mitigates the global gradient distortion seen in vanilla LoRA and better preserves inter-channel balance. Empirically, GraLoRA consistently outperforms standard LoRA and strong baselines such as RaSA across diverse tasks and model scales. On the code generation benchmark HumanEval+, it achieves up to +8.5% absolute gain in Pass1. In commonsense reasoning, GraLoRA delivers improvements across all tasks, with especially strong results on multi-hop and structured reasoning benchmarks.
@@ Appendix C Precise Analysis on Computation Overhead
LoRA FLOPs LoRA performs the projection in two sequential steps to leverage its low-rank structure. In the first step, the computation of $A^{\top}X\in\mathbb{R}^{r\times T}$ requires $(2N-1)rT$ FLOPs. In the second step, the reconstruction $B(A^{\top}X)\in\mathbb{R}^{M\times T}$ incurs $(2r-1)MT$ FLOPs. Therefore, the total FLOPs for LoRA is:
@@
In the first stage (projection), each adapter block computes $A_{i,j}^{\top}X_{j}\in\mathbb{R}^{\frac{r}{k}\times T}$, which requires $\left(2\frac{n}{k}-1\right)\frac{r}{k}T$ FLOPs. Since there are $k^{2}$ such blocks, the total FLOPs for this step is $(2n-k)rT$.
In the second stage (reconstruction), each adapter block performs $B_{i,j}(A_{i,j}^{\top}X_{j})\in\mathbb{R}^{\frac{m}{k}\times T}$, which costs $\left(2\frac{r}{k}-1\right)\frac{m}{k}T$ FLOPs. With $k^{2}$ blocks, the total becomes $(2r-k)mT$.
@@
which requires $\left(\frac{m}{k}\times T\right)(k-1)=\frac{mT(k-1)}{k}$ FLOPs per row. Across $k$ rows, the total cost becomes $(k-1)mT$.
@@
Appendix D Baselines and Hyperparameters
Baseline Methods. We compared GraLoRA with three baseline methods. Key idea for each method is as follows:

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) is a popular method for parameter-efficient fine-tuning (PEFT) of generative models, valued for its simplicity and effectiveness. Despite recent enhancements, LoRA still suffers from a fundamental limitation: overfitting when the bottleneck is widened. It performs best at ranks 32–64, yet its accuracy stagnates or declines at higher ranks, still falling short of full fine-tuning (FFT) performance. We identify the root cause as LoRA’s structural bottleneck, which introduces gradient entanglement to the unrelated input channels and distorts gradient propagation. To address this, we introduce a novel structure, Granular Low-Rank Adaptation (GraLoRA) that partitions weight matrices into sub-blocks, each with its own low-rank adapter. With negligible computational or storage cost, GraLoRA overcomes LoRA’s limitations, effectively increases the representational capacity, and more closely approximates FFT behavior. Experiments on code generation, commonsense reasoning, mathematical reasoning, general language understanding, and image generation benchmarks show that GraLoRA consistently outperforms LoRA and other baselines, achieving up to +8.5% absolute gain in Pass@1 on HumanEval+. These improvements hold across model sizes and rank settings, making GraLoRA a scalable and robust solution for PEFT.
@@ 1 Introduction
Task-specific fine-tuning enables a wide range of applications and significantly improves the quality and effectiveness of generative models. However, the massive scale of these models poses substantial challenges for practical deployment. To address these limitations, Parameter-Efficient Fine-Tuning (PEFT) methods have emerged as a cost-effective alternative Houlsby et al. (2019); Xu et al. (2023). Among them, Low-Rank Adaptation (LoRA) Hu et al. (2021) has gained particular attention for its simplicity and effectiveness, introducing trainable low-rank matrices while keeping the pre-trained model weights frozen. Although the imposed rank-$r$ bottleneck may lead to slight performance degradation compared to full fine-tuning (FFT), its efficiency has led to widespread adoption in practice.
To maximize the benefits of LoRA, various studies have proposed techniques such as improved initialization Büyükakyüz (2024); Meng et al. (2024); Paischer et al. (2024); Wang et al. (2024) and structural refinements He et al. (2025); Huang et al. (2025); Jiang et al. (2024); Kopiczko et al. (2024) to enhance fine-tuning quality. While these efforts have advanced performance, a substantial quality gap remains compared to FFT, largely due to the inherent upper bound on the rank. Although using a higher rank, within hardware limits, appears to be a natural solution, unfortunately, current implementations of LoRA and its variants do not support such flexibility. Simply increasing the rank often leads to degraded accuracy in many scenarios.
@@ 2.1 Introduction to LoRA
$\nabla_{R}L=\frac{\partial L}{\partial R}=\frac{\partial L}{\partial Y}X^{\top}\in\mathbb{R}^{M\times N}$ (2)
@@
These result in the following reconstructed update in the fused weight space:
$\tilde{\nabla}_{R}L=\frac{\partial L}{\partial B}A^{\top}+B\frac{\partial L}{\partial A^{\top}}=\frac{\partial L}{\partial Y}X^{\top}AA^{\top}+BB^{\top}\frac{\partial L}{\partial Y}X^{\top}.$ (4)
@@ 2.2 Why Does LoRA Suffer from a Larger Rank?
These findings reveal a fundamental misalignment between LoRA updates and the gradient landscape shaped by FFT. The entangled influence of input channels caused by the low-rank projection limits LoRA’s ability to selectively learn from salient features, particularly under skewed input statistics. While the negative impact of outliers has been well recognized in the context of quantization Xiao et al. (2024) Lee et al. (2024), their influence on LoRA’s behavior has not been systematically studied until now.
@@ 3.4 Tradeoff Analysis
Selection of k While GraLoRA increases the total rank from $r$ to $kr$, each individual block, represented as $B_{i,j}A_{i,j}^{\top}\in\mathbb{R}^{\frac{M}{k}\times\frac{N}{k}}$, is constrained to a reduced rank of $\frac{r}{k}$. As a result, increasing $k$ beyond a certain threshold can degrade performance due to limited expressiveness within each block. This effect is especially pronounced when the overall rank $r$ is small. Empirically, we observed that maintaining a minimum block expressiveness of approximately $r/k^{2}\approx 8$ yields stable performance across various configurations. Based on this observation, we adopted $k=2$ for ranks 16 and 32, and $k=4$ for ranks 64 and 128 in our experiments. Detailed $k$-sweep results can be found in Section 4.7.
@@ 4 Experiments
In order to validate the superiority of the proposed idea, we conduct an extensive analysis on large-scale dataset with the state-of-the art LLMs. We evaluate GraLoRA across five challenging domains: code generation, commonsense reasoning, mathematical reasoning, general language understanding, and personalized image generation. Our experiments are designed to assess whether the proposed granular adaptation mechanism improves performance across varying model sizes, LoRA ranks, and tasks that require nuanced reasoning and high representational fidelity.
@@ 4.1 Experimental Setup
Code Generation. We fine-tuned LLaMA3.1–8B ( Grattafiori and others (2024)) with 4 A100 80G GPU on the Magicoder-Evol-Instruct-110k Wei et al. (2024) train dataset, a curated and decontaminated subset of WizardCoder Luo et al. (2024), comprising high-quality instruction–response pairs for programming tasks. Evaluation was conducted on the Humaneval+ test dataset following He et al. He et al. (2025), which samples 50 completions per problem using a temperature of 0.2. We report Pass@1, Pass@5, and Pass@10 accuracy following standard protocol via BigCode Evaluation Harness Ben Allal et al. (2022).
Commonsense Reasoning We fine-tuned LLaMA3.2–3B, LLaMA3.1–70B, Qwen-2.5-1.5B, and Qwen-2.5-7B ( Yang et al. (2024)) across 8 commonsense tasks: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-Challenge, ARC-Easy Clark et al. (2018), and OpenBookQA Mihaylov et al. (2018). We followed the training pipeline proposed by LLM-Adapters Hu et al. (2023). Training was performed on 2 H100 80G GPUs for 1.5-8B models, and on 8 A100 80G GPUs for the 70B model. LLaMA3.1–70B, Qwen-2.5-1.5B, and Qwen-2.5-7B were trained with rank 64, using the optimal configurations proposed by Biderman et al. Biderman et al. (2024). LLaMA3.2–3B was trained with rank 32, following the settings of Ponkshe et al. Ponkshe et al. (2025) to ensure a fair comparison with results reported in the original paper.
Mathematical Reasoning We fine-tuned LLaMA3.2–3B on MetaMathQA Yu et al. (2023) train dataset using 4 H100 80G GPUs. Evaluation was done on MATH Hendrycks et al. (2021) dataset, following the evaluation procedure and settings from He et al. He et al. (2025).
General Language Understanding We trained and evaluated RoBERTa-base Liu et al. (2019), an encoder-only architecture model, on the GLUE Wang et al. (2019) benchmark composed of eight sub-tasks. Following the protocol from prior works ( Kopiczko et al. (2024)  Gao et al. (2024)), we excluded MNLI and QQP—two time-intensive tasks—which also meant we did not apply the MNLI-based tricks for MRPC, RTE, and STS-B (as used in the original LoRA paper). Accordingly, we retrained LoRA on these tasks without this optimization and report updated results. All trainings were done on a single H100 80G GPU.
Personalized Image Generation We fine-tuned SDXL Podell et al. (2023) following the official training setup from Huggingface diffusers repository, using the Naruto-Blip-Captions Cervenka (2022) dataset on a single H100 80G GPU. The dataset was split 90% for training and 10% for evaluation. The quality was measured through CLIP similarity and DINOv2 similarity scores.
Training Details We conducted experiments on five open-sourced LLMs—LLaMA3.1–8B, LLaMA3.1–70B, LLaMA3.2–3B, Qwen-2.5-1.5B, and Qwen-2.5-7B—covering diverse architecture and sclaes across code generation, commonsense reasoning, and mathematical reasoning tasks. Following common practice ( Jiang et al. (2024); Kopiczko et al. (2024)), we used pre-trained models rather than instruction-tuned models. All PEFT methods were applied to the linear modules in both the attention ( $W_{q},W_{k},W_{v},W_{o}$ )and the feed-forward networks ($W_{up},W_{down},W_{gate}$). We adopted alpaca-chat instruction template for training and evaluation. We compared GraLoRA to three representative PEFT methods: LoRA, MoRA  Jiang et al. (2024) and RaSA  He et al. (2025). We have also handled RoBERTa-base and SDXL, to show the robustness and scalability of our method across differnt models and tasks. Hyperparameters for GraLoRA followed those introduced in Kopiczko et al. Kopiczko et al. (2024), except for learning rate, which was reduced by a factor of 5–10, as VeRA uses a learning rate approximately 10 times larger than LoRA. Detailed training parameters can be found in Appendix E.
@@ 4.2 Results on Code Generation
[…]
@@ 4.3 Results on Commonsense Reasoning
Table 3: Commonsense reasoning accuracy across models and tasks. Bold indicates the best performance per column. HS means HellaSwag, and WG WinoGrande. indicates values reported by Ponkshe et al. Ponkshe et al. (2025)
@@
As shown in Table 3, GraLoRA outperformed other methods across a wide range of models and tasks. Notably, GraLoRA demonstrated superior performance across models of varying scales, achieving a 1.1% improvement in average accuracy on both Qwen2.5-1.5B and LLaMA3.1-70B. It also yielded a 0.9% gain on the widely used mid-sized model, Qwen2.5-7B. Moreover, GraLoRA achieved a 3.3% improvement on LLaMA3.2-3B, surpassing a broad range of baselines as presented in Table 6.
Furthermore, GraLoRA achieved the best results on 26 out of 32 tasks, consistently outperforming alternatives across benchmarks. These results support our analysis in Section 3.3, showing that GraLoRA’s localized updates enhance alignment with FFT and promote robust generalization in multi-aspect reasoning tasks.
4.4 Results on Mathematical Reasoning
Table 4: MATH dataset accuracy results on Qwen2.5-1.5B using LoRA and GraLoRA across different ranks. Best results per group are in bold.
Rank Method Training Time Relative Time Accuracy
[…]
In mathematical reasoning task, regarded as one of the most challenging benchmarks, GraLoRA consistently outperformed LoRA across all configurations. Notably, in the high rank setting of $r=128$, GraLoRA achieved a 4.2% improvement in accuracy (Table 4), mirroring the performance trends observed in the code generation experiments. These results further highlight the robustness of GraLoRA, demonstrating its capability to fully exploit the advantages enabled by increased rank capacity, thereby overcoming the inherent expressiveness constraints of previous PEFT methods.
4.5 Results on General Language Understanding
GraLoRA demonstrates strong performance even in the low-rank regime, outperforming all baselines in terms of average score. The Hybrid GraLoRA variant achieves the most robust results, attaining the best performance on four out of six tasks, while both the original and hybrid versions consistently surpass all other baselines, as shown in Table 7. Compared with LoRA, the best GraLoRA configuration yields a 1.8% improvement in average accuracy, with gains observed across all sub-tasks. These findings indicate that GraLoRA maintains high effectiveness even under constrained parameter budgets and generalizes well to non-LLM architectures.
4.6 Results on Personalized Image Generation
In the image generation task, GraLoRA consistently outperformed LoRA in both CLIP and DINOv2 similarity metrics, achieving 0.5% and 2.1% improvements, respectively (Table 8). These results further demonstrate the generality and effectiveness of GraLoRA beyond language models, extending its applicability to vision–language and generative architectures such as diffusion models.
@@ 4.7 Ablation Study
Figure 8: (a) GraLoRA $k$ sweep results and (b) Hybrid GraLoRA Ratio sweep results for LLaMA3.1–8B on code generation task. Ratio 0 implies default GraLoRA and ratio 1 vanilla LoRA in (b).
@@ 5 Conclusion
GraLoRA addresses this limitation by dividing the adaptation space into $k^{2}$ independently trained low-rank adapters, enabling spatially localized and context-aware updates. Our theoretical analysis shows that this design increases expressivity by a factor of $k$, without additional parameters or computational cost. Moreover, under outlier activations, GraLoRA effectively mitigates the global gradient distortion seen in vanilla LoRA and better preserves inter-channel balance. Empirically, GraLoRA consistently outperforms standard LoRA and strong baselines such as RaSA across diverse tasks and model scales. On the code generation benchmark HumanEval+, it achieves up to +8.5% absolute gain in Pass1. GraLoRA also delivers significant improvements across other 4 additional tasks, highlighting its robustness and scalability across heterogeneous architectures and model sizes.
@@
Acknowledgments and Disclosure of Funding
This work was partly supported by the Technology development Program of MSS [RS-2023-00258531], the Starting growth Technological R&D Program of MSS [S3358777], and the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2025-02304183, Development of Optimization Code Conversion Technology for Heterogeneous Al Semiconductor-Based Large-Scale Models).
@@ Appendix C Precise Analysis on Computation Overhead
LoRA FLOPs LoRA performs the projection in two sequential steps to leverage its low-rank structure. In the first step, the computation of $A^{\top}X\in\mathbb{R}^{r\times T}$ requires (2N1)rT(2N-1)rT FLOPs. In the second step, the reconstruction $B(A^{\top}X)\in\mathbb{R}^{M\times T}$ incurs (2r1)MT(2r-1)MT FLOPs. Therefore, the total FLOPs for LoRA is:
@@
In the first stage (projection), each adapter block computes $A_{i,j}^{\top}X_{j}\in\mathbb{R}^{\frac{r}{k}\times T}$, which requires $\left(2\frac{n}{k}-1\right)\frac{r}{k}T$ FLOPs. Since there are $k^{2}$ such blocks, the total FLOPs for this step is (2nk)rT(2n-k)rT.
In the second stage (reconstruction), each adapter block performs $B_{i,j}(A_{i,j}^{\top}X_{j})\in\mathbb{R}^{\frac{m}{k}\times T}$, which costs $\left(2\frac{r}{k}-1\right)\frac{m}{k}T$ FLOPs. With $k^{2}$ blocks, the total becomes (2rk)mT(2r-k)mT.
@@
which requires $\left(\frac{m}{k}\times T\right)(k-1)=\frac{mT(k-1)}{k}$ FLOPs per row. Across $k$ rows, the total cost becomes (k1)mT(k-1)mT.
@@
Appendix D Additional Experiment Results
D.1 Results on Commonsense Reasoning with extensive baseline comparison
Table 6: Commonsense reasoning accuracy across models and tasks. All values are percentages; bold indicates the best performance per row. HS means HellaSwag, and WG WinoGrande. indicates values reported by Ponkshe et al. [25]
Method Rank #Parmas BoolQ PIQA SIQA HS WG ARC-c ARC-e OBQA Avg.
Full-FT - 3.21B 70.4% 85.6% 80.5% 91.9% 85.0% 75.3% 88.5% 81.9% 82.4%
LoRA-XS 96 1.81M 67.3% 83.4% 78.7% 89.0% 82.1% 72.6% 85.2% 78.9% 79.6%
LoRA-SB 96 1.81M 70.3% 84.8% 80.2% 91.6% 84.6% 74.7% 87.9% 81.2% 81.9%
LoRA 32 48.63M 70.0% 85.2% 79.1% 90.7% 82.2% 74.3% 86.9% 81.9% 81.3%
MELoRA 32 48.63M 71.3% 85.0% 78.6% 93.0% 79.7% 73.7% 85.5% 79.0% 80.7%
rsLoRA 32 48.63M 69.8% 85.1% 78.9% 90.5% 82.0% 74.2% 86.7% 81.7% 81.1%
PiSSA 32 48.63M 70.1% 85.4% 79.4% 90.9% 82.7% 74.6% 87.2% 81.8% 81.5%
DoRA 32 49.40M 70.4% 85.6% 79.7% 90.8% 82.9% 74.9% 87.6% 82.0% 81.7%
BOFT 32 48.48M 72.3% 84.6% 79.1% 91.3% 84.5% 73.7% 87.8% 80.6% 81.7%
LoRA-Pro 32 48.63M 71.3% 85.8% 79.4% 90.9% 83.4% 75.3% 87.2% 81.7% 81.9%
MoRA 32 48.63M 72.4% 86.1% 80.1% 92.3% 84.8% 76.8% 88.8% 84.8% 83.3%
RaSA 32 48.63M 73.1% 87.5% 81.1% 93.7% 85.3% 78.9% 88.9% 83.6% 84.0%
GraLoRA 32 48.63M 74.1% 86.5% 80.8% 93.8% 87.5% 79.9% 89.5% 84.8% 84.6%
D.2 Results on General Language Understanding (GLUE)
Table 7: GLUE dataset accuracy results on RoBERTa-base across different tasks. Best results per task are in bold.
Method #Params SST-2 (%) MRPC (%) CoLA (%) QNLI (%) RTE (%) STS-B (%) Avg (%)
Full-FT 125M 94.8 90.2 63.6 92.8 78.7 91.2 85.2
LoRA 0.3M 95.1 86.5 63.4 93.3 76.2 90.6 84.2
VeRA 0.043M 94.6 89.5 65.6 91.8 78.7 90.7 85.2
FourierFT 0.024M 94.2 90.0 63.8 92.2 79.1 90.8 85.0
GraLoRA 0.3M 95.2 89.7 65.3 93.0 80.9 91.1 85.8
Hybrid GraLoRA 0.3M 95.2 90.2 64.1 93.4 79.8 91.2 85.6
Best GraLoRA 0.3M 95.2 90.2 65.3 93.4 80.9 91.2 86.0
D.3 Results on Personalized Image Generation
Table 8: SDXL fine-tuning results personalized image generation.
Method CLIP Similarity DINOv2 Similarity
LoRA 91.4% 79.2%
GraLoRA 91.9% 81.3%
Appendix E Experiment Details
Baseline Methods. We compared GraLoRA with three main baseline methods. Key idea for each method is as follows: